Task
1.) 32^ 2/5
2.)125 2/3
3)8 2/3
4.)16 2/3
5.)4 3/2
6.)32 2/5+64 2/3
1
Expert's answer
2012-08-28T08:22:38-0400
1.) 32^ 2/5=(2^5)^2/5=2^2=4
& 2.)125^2/3=(5^3)^2/3=25
& 6.)32^2/5+64^2/3=(2^5)^2/5+(2^6)^2/3=4+16=20
and so on